Please help with all parts below:

Problem 15.76 The contribution of the combustion of various fuels to the buildup of CO2 in the atmosphere can be assessed in different ways. One way relates the mass of CO2 formed to the mass of fuel burned; another relates the mass of CO2 to the quantity of heat evolved in the combustion. The heat released per mole of each substance is C (graphite), 393.5 kJ C2Ho (g), 1554 kJ, and C4H10 (g), 2877 kJ Which of the three fuels, C(graphite), C2Ho(g), or C H10 (g) produces the smallest mass of CO2 per gram of fuel? C(graphite) C2H6(g) O C H10 (g) Which of the three fuels, C(graphite), C2H6(g), or C H10 (g), produces the smallest mass of CO2 per gram per kilojoule of heat evolved? C(graphite) C2H6(g) H10 (g)
Answer #1

Ans to problem 15.76

1.) C2H6 (g)

Mass of CO2 produced per gram of fuel = mass of CO2 produced/ molar mass of fuel

graphite = 44/12 = 3.67g

For C2H6 = 88/30 = 2.93g

for C4H10 = 176/58 = 3.03g

2.) C4H10 (g)

Amount of CO2 per gram per kilojoule = mass of CO2 formed divided by molecular mass divided by heat released in kilojoule

for graphite = 3.66/ 393.5 = 0.0093 g

for ethane = 2.933/1554 = 0.0018g

for butane = 3.03/ 2877 = 0.0010 g

so smallest is for butane